# Score-based Generative Diffusion Models for Social Recommendations

### v1

@@ Abstract
With the prevalence of social networks on online platforms, social recommendation has become a vital technique for enhancing personalized recommendations. The effectiveness of social recommendations largely relies on the social homophily assumption, which presumes that individuals with social connections often share similar preferences. However, this foundational premise has been recently challenged due to the inherent complexity and noise present in real-world social networks. In this paper, we tackle the low social homophily challenge from an innovative generative perspective, directly generating optimal user social representations that maximize consistency with collaborative signals. Specifically, we propose the Score-based Generative Model for Social Recommendation (SGSR), which effectively adapts the Stochastic Differential Equation (SDE)-based diffusion models for social recommendations. To better fit the recommendation context, SGSR employs a joint curriculum training strategy to mitigate challenges related to missing supervision signals and leverages self-supervised learning techniques to align knowledge across social and collaborative domains. Extensive experiments on real-world datasets demonstrate the effectiveness of our approach in filtering redundant social information and improving recommendation performance. Our codes are available at  https://github.com/Anonymous-CodeRepository/Score-based-Generative-Diffusion-Models-for-Social-Recommendations-SGSR
@@
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
In the era of information explosion, Recommender Systems (RS) have emerged as indispensable tools for mitigating information overload and personalizing users’ content stream across e-commerce and social media platforms. Collaborative Filtering (CF), which aims to learn user and item representations from historical interactions, forms the foundation of modern recommender systems  [1, 2, 3, 4]. However, CF-based methods face a persistent challenge due to their reliance on the quality of user-item interaction data: real-world interaction data used for training is often sparse, which can adversely impact the performance of these models  [5, 6, 7]. To address the data sparseness issue, social recommendation methods incorporate social context to enhance the modelling of user intent within user representations. This practice aligns with the idea that individuals connected through social ties often exhibit similar preferences—a phenomenon supported by social influence theory [8, 9] and also recognized as social homophily  [10, 11]. In recent years, a wide range of research has focused on enhancing the performance of social recommendations by leveraging the social homophily condition, encompassing Graph Neural Networks (GNNs) methods [12, 13, 14, 15] and Self-Supervised Learning (SSL) methods [16, 17].
@@
Nevertheless, the underlying assumption of social homophily has recently been questioned. Extensive statistical and empirical analyses  [18, 19, 11, 20] have suggested that leveraging the raw social graph can be unreliable and may further degrade the recommendation performance. This counter-intuition arises because of the inherent complexity and noise present in real-world social networks  [21, 22], introducing noisy information into learned user representations. Considering the example in Fig. 1, a male student (i.e., User A) might establish social relationships with a female classmate (i.e., User B) who has recently acquired cosmetics or feminine apparel, yet such relationships likely contribute to unreliable recommendations for the male. Consequently, these noisy social signals can propagate through social encoder layers, accumulating bias and ultimately hindering the recommender system’s effectiveness  [19]. Thus, it is necessary to develop social denoising methods to filter redundant social signals from user-user social graph, alleviating the negative impact of such signals on recommendation performance.
@@
  • Expressive Capability Limitation: These methods are unable to adaptively account for all users within a social network, which is computationally infeasible due to the vast number of users. The graph rewiring algorithms focus on a small, predefined subset with a fixed number of candidates, such as one-hop neighbours or users with high embedding similarity. This approach neglects users with similar preferences who do not conform to these predefined criteria, thereby missing opportunities to enhance the alignment between social and collaborative signals. Furthermore, extending denoising to multi-hop optimization may introduce conflicts between local and global optima in the graph editing process.
@@
To overcome the aforementioned limitations, we propose a fundamentally novel approach to denoise social information in recommendation systems (RS) from a generative perspective. Specifically, our key idea is to model the underlying distribution of users’ social representations and directly generate denoised social representations that maximize consistency with collaborative signals. This generative paradigm is highly desirable as it allows us to transcend all possible graph structure revisions and achieve the ideal social representation, which may be unreachable through graph editing steps under computational constraints. To instantiate this novel generative paradigm, we draw inspiration from the success of diffusion models in image restoration [24, 25, 26], and propose to leverage the potential of score-based generative models (SGMs) [27], a strong yet unexplored class of diffusion models. Empowered by stochastic differential equations (SDEs), SGMs unify the diffusion processes, such as DDPM [28] and SMLD [29], in continuous time steps. This unified continuous formulation enables a predict-and-correct generative paradigm that flexibly incorporates a wide range of numerical SDE solvers and statistical methods, thereby reducing numerical errors in the computation of diffusion models and more accurately capturing the complexities of the social denoising problem [30, 31].
@@
Despite the great potential, it is highly challenging to directly take advantage of SGMs for social denoising in social recommendations. First, the optimal social network that accurately aligns with users’ preference patterns remains unknown, leading to a lack of appropriate supervision signals (i.e., training objectives) for SGMs. Additionally, since SGMs were originally designed for image generation and denoising, applying them to social recommendations introduces multiple challenges, such as handling high-order graph structures and aligning signals from social and collaborative domains. To address the challenges above, we propose the Score-based Generative Model for Social Recommendation (SGSR) to transcend the social homophily condition in social recommendations with SGMs. Instead of manually determining the sub-optimal embeddings from edited social graphs as supervisory signals, a curriculum learning mechanism is introduced to facilitate the modelling of social representations across social graphs with incrementally sparse levels. Furthermore, a joint optimization method is developed to leverage collaborative signals to implicitly identify the optimal social representation. Our main contributions are summarized as follows:
@@ II-C Score-based Generative Social Denoising
Additionally, we obtain user and item representations from the collaborative domain to control the social denoising process. We adopt an arbitrary graph-based encoder (e.g., LightGCN [3], PinSAGE [35], etc.) for collaborative filtering with $L$ layers, denoted by $\text{{RecEnc}}(\cdot)$, and obtain the collaborative representation as follows:
@@
Forward Process. The forward process aims to transform the original data distribution of social representations into a prior distribution (e.g., Gaussian noise) in a continuous manner. Considering the forward diffusion process ${\{\mathbf{x}_{t}\}}_{t\in[0,T]}$ with respect to the continuous time step $t$, we first define the raw social embedding $(\mathbf{e}_{S}^{(L)})_{u}$ of a user $u$ as the initial state $\mathbf{x}_{0}$, taken from embedding matrix $\mathbf{E}_{S}^{(L)}$, and $\mathbf{x_{T}}$ follows the prior distribution. The forward diffusion process is defined by the following SDE:
@@
where $\mathbf{w}$ refers to the standard Wiener process [36] (a.k.a. Brownian motion), and $\mathrm{d}t$ is an infinitesimal time step. The vector-valued drift coefficient $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ denotes the drift coefficient determining the data perturbation direction, and the scalar diffusion coefficient $g:\mathbb{R}\rightarrow\mathbb{R}$ represents time-dependent noise magnitude. The drift coefficient and scalar function can be manually defined to effectively guide the perturbation toward the known prior.
Reverse Process. The reverse-time SDE simulates the diffusion process in reverse time, reconstructing samples from a known prior that adhere to the original data distribution. We run the reverse process conditioned on the user’s collaborative information to guide the recovery of the user representation towards a collaboratively consistent direction. This can be formulated as the following conditioned reverse-time SDE [7]:
@@
$\min_{\boldsymbol{\theta}}\ \mathbb{E}_{t}\left[\lambda(t)\mathbb{E}_{\mathbf{x}_{0}}\mathbb{E}_{\mathbf{x}_{t}|\mathbf{c}}\left\lVert{\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{c})\!-\!{\nabla}_{\mathbf{x}_{t}}{\log p_{t}(\mathbf{x}_{t}|\mathbf{c})}}\right\rVert^{2}\right],$ (5)
@@
$\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t}|\mathbf{c})=\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t})+\nabla_{\mathbf{x}}\log p(\mathbf{\mathbf{c}|\mathbf{x}_{t}}),$ (6)
@@
$\ \mathbb{E}_{t}[\lambda(t)\mathbb{E}_{\mathbf{x}_{0}}\mathbb{E}_{\mathbf{x}_{t}|\mathbf{c}}\left\lVert{\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{c})\!-\!{\nabla}_{\mathbf{x}_{t}}{\log p_{t}(\mathbf{x}_{t}|\mathbf{c})}}\right\rVert^{2}]$ (7)
$=\!\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{c}}[\lambda(t){\parallel\!\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{c})\!-\!\nabla\log p_{t}(\mathbf{x_{t}}|\mathbf{c})\!\parallel}^{2}]$
$=\!\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{c}}[\lambda(t){\parallel\!\mathbf{s}_{\boldsymbol{\theta}}(x_{t},t,\mathbf{c})\!-\!\nabla\log p_{t}(\mathbf{x_{t}},\mathbf{c})\!+\!\nabla\log p(\mathbf{c})\!\parallel}^{2}]$
$=\!\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{c}}[\lambda(t){\parallel\!\mathbf{s}_{\boldsymbol{\theta}}(x_{t},t,\mathbf{c})\!\parallel}^{2}\!-\!2\!\cdot\!\mathbf{s}_{\boldsymbol{\theta}}(x_{t},t,\mathbf{c})\!\cdot\!\nabla\log p_{0}t(\mathbf{x}_{t}|\mathbf{x}_{0})]\!+\!C_{1}$
$=\!\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{c}}[\lambda(t){\parallel\!\mathbf{s}_{\boldsymbol{\theta}}(x_{t},t,\mathbf{c})\!-\!\nabla\log p_{0t}(\mathbf{x_{t}}|\mathbf{x_{0}})\parallel}^{2}]\!-\!C_{2}\!+\!C_{1}$
$=\!\mathbb{E}_{t}[\lambda(t)\mathbb{E}_{\mathbf{x}_{0}}\mathbb{E}_{\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{c}}\left\lVert{\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{c})\!-\!{\nabla}_{\mathbf{x}_{t}}{\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})}}\right\rVert^{2}]\!-\!C_{2}\!+\!C_{1}.$
Thus, the denoising score function ${\nabla}_{\mathbf{x}_{t}}{\log p_{t}(\mathbf{x}_{t})}$ depends only on the noising process and can be evaluated in closed form. Finally, the SGSR employs the following classifier-free objectives to optimize the conditional scoring network $s_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$:
$\min_{\boldsymbol{\theta}}\mathbb{E}_{t}\!\left[\lambda(t)\mathbb{E}_{\mathbf{x}_{0}}\mathbb{E}_{\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{c}}\left\lVert{\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{c})\!-\!{\nabla}_{\mathbf{x}_{t}}{\log p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})}}\right\rVert^{2}\right]\!,\!$ (8)
where $\lambda(t)$ is empirically set to 1 based on experimental results.
SDE Instantiation. Despite the universality of the score-based generative paradigm, it is essential to adopt specific designs for the drift coefficient $\mathbf{f}(\boldsymbol{x},t)$ and diffusion coefficient $g(t)$ to instantiate the SGM into practice. to instantiate the SGM in practice. To this end, we implement the drift and diffusion coefficients using two widely adopted formulations: the Variance Exploding (VE) SDE, a continuous form of the Score Matching with Langevin Dynamics (SMLD) [29] model, and the Variance Preserving (VP) SDE, a continuous form of the well-known Denoising Diffusion Probabilistic Models [28]. Specifically, the forward processes of these two SDEs are formalized as follows:
@@
where $\sigma(t)$ and $\beta(t)$ represent the predefined noise scalers. The reverse processes of these SDEs can be obtained via Eq. (4).
@@
[…]
Thus, with the closed-form Gaussian transition kernel, it is straightforward to derive the corresponding probability density function and obtain the closed-form solution of the score functions.
Social Reperesentation Denoising. After training the score-based social denoising model, we obtain the scoring network $\mathbf{s}_{\boldsymbol{\theta}}$ and use it to denoise users’ social representations. Given a user’s raw social representation $(\mathbf{e}_{S}^{(L)})_{u}$, the model first perturbs the raw social representation in a continuous manner using either the forward VP SDE or VE SDE in Eq. (10) to obtain the output $\mathbf{x}_{T}$. It then reconstructs the denoised representation $(\mathbf{\hat{e}}_{S}^{(L)})_{u}$ by iteratively solving the collaboratively conditioned reverse-SDE (Eq. 4) with off-the-shelf SDE initial-value problem (IVP) numerical solvers [38]. To obtain the numerical solution of a given SDE, we adopt the Predictor-Corrector (PC) method [27], a flexible and powerful framework for solving SDEs inherent in SGMs. The iterative steps of the Predictor-Corrector solver first use the prediction of numerical SDE solvers like the Euler–Maruyama method [39], and then correct the predicted trajectory with score-based MCMC methods, such as Langevin MCMC [40], achieving a balance of numerical accuracy and efficiency. The Algorithm  1 outlines the inference process of the denoising model, incorporating annealed Langevin dynamics as a corrector ($\epsilon_{i}$ represents the step size for Langevin Dynamics).

### v2

@@ Abstract
With the prevalence of social networks on online platforms, social recommendation has become a vital technique for enhancing personalized recommendations. The effectiveness of social recommendations largely relies on the social homophily assumption, which presumes that individuals with social connections often share similar preferences. However, this foundational premise has been recently challenged due to the inherent complexity and noise present in real-world social networks. In this paper, we tackle the low social homophily challenge from an innovative generative perspective, directly generating optimal user social representations that maximize consistency with collaborative signals. Specifically, we propose the Score-based Generative Model for Social Recommendation (SGSR), which effectively adapts the Stochastic Differential Equation (SDE)-based diffusion models for social recommendations. To better fit the recommendation context, SGSR employs a joint curriculum training strategy to mitigate challenges related to missing supervision signals and leverages self-supervised learning techniques to align knowledge across social and collaborative domains. Extensive experiments on real-world datasets demonstrate the effectiveness of our approach in filtering redundant social information and improving recommendation performance.
@@
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ I Introduction
In the era of information explosion, Recommender Systems (RS) have emerged as indispensable tools for mitigating information overload and personalizing users’ content stream across e-commerce and social media platforms. Collaborative Filtering (CF), which aims to learn user and item representations from historical interactions, forms the foundation of modern recommender systems  [1, 2, 3, 4]. However, CF-based methods face a persistent challenge due to their reliance on the quality of user-item interaction data: real-world interaction data used for training is often sparse, which can adversely impact the performance of these models  [5, 6, 7]. To address the data sparseness issue, social recommendation methods incorporate social context to enhance the modelling of user intent within user representations. This practice aligns with the idea that individuals connected through social ties often exhibit similar preferences—a phenomenon supported by social influence theory [8] and also recognized as social homophily  [9, 10]. In recent years, a wide range of research has focused on enhancing the performance of social recommendations by leveraging the social homophily condition, encompassing Graph Neural Networks (GNNs) methods [11, 12, 13, 14] and Self-Supervised Learning (SSL) methods [15, 16].
@@
Nevertheless, the underlying assumption of social homophily has recently been questioned. Extensive statistical and empirical analyses  [17, 18, 10, 19] have suggested that leveraging the raw social graph can be unreliable and may further degrade the recommendation performance. This counter-intuition arises because of the inherent complexity and noise present in real-world social networks  [20], introducing noisy information into learned user representations. Considering the example in Fig. 1, a male student (i.e., User A) might establish social relationships with a female classmate (i.e., User B) who has recently acquired cosmetics or feminine apparel, yet such relationships likely contribute to unreliable recommendations for the male. Consequently, these noisy social signals can propagate through social encoder layers, accumulating bias and ultimately hindering the recommender system’s effectiveness  [18]. Thus, it is necessary to develop social denoising methods to filter redundant social signals from user-user social graph, alleviating the negative impact of such signals on recommendation performance.
@@
  • Expressive Capability Limitation: These methods are unable to adaptively account for all users within a social network, which is computationally infeasible due to the vast number of users. The graph rewiring algorithms focus on a small, predefined subset with a fixed number of candidates, such as one-hop neighbours or users with high embedding similarity. This approach neglects users with similar preferences who do not conform to these predefined criteria, thereby missing opportunities to enhance the alignment between social and collaborative signals. Furthermore, extending denoising to multi-hop optimization may introduce conflicts between local and global optima in the graph editing process. This suggests that better social connections may not necessarily yield optimized social representations, compounded by the inherent noise in message passing approaches as validated by previous works [18].
@@
To overcome the aforementioned limitations, we propose a fundamentally novel approach to denoise social information in recommendation systems (RS) from a generative perspective. Specifically, our key idea is to model the underlying distribution of users’ social representations and directly generate denoised social representations that maximize consistency with collaborative signals. This generative paradigm is highly desirable as it allows us to transcend all possible graph structure revisions and achieve the ideal social representation, which may be unreachable through graph editing steps under computational constraints. To instantiate this novel generative paradigm, we draw inspiration from the success of diffusion models in image restoration [22], and propose to leverage the potential of score-based generative models (SGMs) [23], a strong yet unexplored class of diffusion models. Empowered by stochastic differential equations (SDEs), SGMs unify the diffusion processes, such as DDPM [24] and SMLD [25], in continuous time steps. This unified continuous formulation enables a predict-and-correct generative paradigm that flexibly incorporates a wide range of numerical SDE solvers and statistical methods, thereby reducing numerical errors in the computation of diffusion models and more accurately capturing the complexities of the social denoising problem [26, 27].
@@
Despite the great potential, it is highly challenging to directly take advantage of SGMs for social denoising in social recommendations. First, the optimal social network that accurately aligns with users’ preference patterns remains unknown, leading to a lack of appropriate supervision signals (i.e., training objectives) for SGMs. Additionally, while SGMs were originally developed for image generation/denoising, their direct application faces two fundamental difficulties in social recommendation: (1) effectively handling high-order graph structures, and (2) aligning signals across social and collaborative domains. To address the challenges above, we propose the Score-based Generative Model for Social Recommendation (SGSR) to transcend the social homophily condition in social recommendations with SGMs. Instead of manually determining the sub-optimal embeddings from edited social graphs as supervisory signals, a curriculum learning mechanism is introduced to facilitate the modelling of social representations across social graphs with incrementally sparse levels. Furthermore, a joint optimization method is developed to leverage collaborative signals to implicitly identify the optimal social representation. Our main contributions are summarized as follows:
@@ II-C Score-based Generative Social Denoising
Additionally, we obtain user and item representations from the collaborative domain to control the social denoising process. We adopt an arbitrary graph-based encoder (e.g., LightGCN [2], etc.) for collaborative filtering with $L$ layers, denoted by $\text{{RecEnc}}(\cdot)$, and obtain the collaborative representation as follows:
@@
Forward Process. The forward process aims to transform the original data distribution of social representations into a prior distribution (e.g., Gaussian noise) in a continuous manner. Considering the forward diffusion process ${\{\mathbf{x}_{t}\}}_{t\in[0,T]}$ with respect to the continuous time step $t$, we first define the raw social embedding $(\mathbf{e}_{S}^{(L)})_{u}$ of a user $u$ as the initial state $\mathbf{x}_{0}$, taken from embedding matrix $\mathbf{E}_{S}^{(L)}$, and $\mathbf{x_{T}}$ follows a standard Gaussian prior distribution. The forward diffusion process is defined by the following SDE:
@@
where $\mathbf{w}$ refers to the standard Wiener process (a.k.a. Brownian motion), and $\mathrm{d}t$ is an infinitesimal time step. The vector-valued drift coefficient $f:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ denotes the drift coefficient determining the data perturbation direction, and the scalar diffusion coefficient $g:\mathbb{R}\rightarrow\mathbb{R}$ represents time-dependent noise magnitude. The drift coefficient and scalar function can be manually defined to effectively guide the perturbation toward the known prior.
Reverse Process. The reverse-time SDE simulates the diffusion process in reverse time, reconstructing samples from a known prior that adheres to the original data distribution, which aligns with the denoising objective. We run the reverse process conditioned on the user’s collaborative information to guide the recovery of the user representation towards a collaboratively consistent direction. This can be formulated as the following conditioned reverse-time SDE [30]:
@@
$\hskip-5.97508pt\min_{\boldsymbol{\theta}}\ \mathbb{E}_{t}\left[\lambda(t)\mathbb{E}_{\mathbf{x}_{0}}\mathbb{E}_{\mathbf{x}_{t}|\mathbf{c}}\left\lVert{\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{c})\!-\!{\nabla}_{\mathbf{x}_{t}}{\log p_{t}(\mathbf{x}_{t}|\mathbf{c})}}\right\rVert^{2}\right],$ (5)
@@
$\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t}|\mathbf{c})=\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x}_{t})+\nabla_{\mathbf{x}}\log p_{t}(\mathbf{\mathbf{c}|\mathbf{x}_{t}}),$ (6)
@@
$\hskip-7.68222pt\begin{aligned} &\ \mathbb{E}_{t}[\lambda(t)\mathbb{E}_{\mathbf{x}_{0}}\mathbb{E}_{\mathbf{x}_{t}|\mathbf{c}}\left\lVert{\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{c})\!-\!{\nabla}_{\mathbf{x}_{t}}{\log p_{t}(\mathbf{x}_{t}|\mathbf{c})}}\right\rVert^{2}]\\ &=\!\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{c}}[\lambda(t){\parallel\!\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{c})\!-\!\nabla\log p_{t}(\mathbf{x_{t}}|\mathbf{c})\!\parallel}^{2}]\\ &=\!\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{c}}[\lambda(t){\parallel\!\mathbf{s}_{\boldsymbol{\theta}}(x_{t},t,\mathbf{c})\!-\!\nabla\log p_{t}(\mathbf{x_{t}},\mathbf{c})\!+\!\nabla\log p(\mathbf{c})\!\parallel}^{2}]\\ &=\!\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{c}}[\lambda(t){\parallel\!\mathbf{s}_{\boldsymbol{\theta}}(x_{t},t,\mathbf{c})\!\parallel}^{2}\!-\!2\!\cdot\!\mathbf{s}_{\boldsymbol{\theta}}(x_{t},t,\mathbf{c})\!\cdot\!\nabla\log p_{0}t(\mathbf{x}_{t}|\mathbf{x}_{0})]\!+\!C_{1}\\ &=\!\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t},\mathbf{c}}[\lambda(t){\parallel\!\mathbf{s}_{\boldsymbol{\theta}}(x_{t},t,\mathbf{c})\!-\!\nabla\log p_{0t}(\mathbf{x_{t}}|\mathbf{x_{0}})\parallel}^{2}]\!-\!C_{2}\!+\!C_{1}\\ &=\!\mathbb{E}_{t}[\lambda(t)\mathbb{E}_{\mathbf{x}_{0}}\mathbb{E}_{\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{c}}\left\lVert{\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{c})\!-\!{\nabla}_{\mathbf{x}_{t}}{\log p_{0t}(\mathbf{x}_{t}|\mathbf{x}_{0})}}\right\rVert^{2}]\!-\!C_{2}\!+\!C_{1},\end{aligned}$ (7)
where $p_{0t}(\mathbf{x}_{t}|\mathbf{x}_{0})$ represents the transition kernal from $\mathbf{x}_{0}$ to $\mathbf{x}_{t}$. Thus, the denoising score function ${\nabla}_{\mathbf{x}_{t}}{\log p_{t}(\mathbf{x}_{t})}$ depends only on the noising process and can be evaluated in closed form. Finally, the SGSR employs the following classifier-free objectives to optimize the conditional scoring network $s_{\theta}(\mathbf{x}_{t},t,\mathbf{c})$:
$\hskip-5.69054pt\min_{\boldsymbol{\theta}}\mathbb{E}_{t}\!\left[\lambda(t)\mathbb{E}_{\mathbf{x}_{0}}\mathbb{E}_{\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{c}}\left\lVert{\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\mathbf{c})\!-\!{\nabla}_{\mathbf{x}_{t}}{\log p_{0t}(\mathbf{x}_{t}|\mathbf{x}_{0})}}\right\rVert^{2}\right]\!,\!$ (8)
where $\lambda(t)$ is empirically set to 1 based on experiments.
SDE Instantiation. Despite the universality of the score-based generative paradigm, it is essential to adopt specific designs for the drift coefficient $\mathbf{f}(\boldsymbol{x},t)$ and diffusion coefficient $g(t)$ to instantiate the SGM into practice. To this end, we implement the drift and diffusion coefficients using two widely adopted formulations: the Variance Exploding (VE) SDE, a continuous form of the Score Matching with Langevin Dynamics (SMLD) [25] model, and the Variance Preserving (VP) SDE, a continuous form of the well-known Denoising Diffusion Probabilistic Models [24]. Specifically, the forward processes of these two SDEs are formalized as follows:
@@
where $\sigma(t)$ and $\beta(t)$ represent the predefined noise scalers. The reverse processes of the SDEs can be obtained via Eq. (4).
@@
$p_{t}(\mathbf{x}_{t}|\mathbf{x}_{0})\!=\!\begin{cases}\!\mathcal{N}\left(\mathbf{x}_{t};\mathbf{x}_{0},[{\sigma}^{2}(t)-{\sigma}^{2}(0)]\mathbf{I}\right),\!&(\text{VE SDE})\\ \!\mathcal{N}(\mathbf{x}_{t};\mathbf{x}_{0}{e}^{-\frac{1}{2}\int^{t}_{0}\beta(s)ds},\mathbf{I}-\mathbf{I}{e}^{-\int^{t}_{0}\beta(s)ds}),\!&(\text{VP SDE})\end{cases}$ (10)
The closed-form Gaussian transition kernel enables the direct derivation of both the probability density function and the analytical solutions of the score functions.
Social Reperesentation Denoising. After training the score-based social denoising model, we obtain the scoring network $\mathbf{s}_{\boldsymbol{\theta}}$ and use it to denoise users’ social representations. Given a user’s raw social representation $(\mathbf{e}_{S}^{(L)})_{u}$, the model first perturbs the raw social representation in a continuous manner using either the forward VP SDE or VE SDE in Eq. (9) to obtain the output $\mathbf{x}_{T}$. It then reconstructs the denoised representation $(\mathbf{\hat{e}}_{S}^{(L)})_{u}$ by iteratively solving the collaboratively conditioned reverse-SDE (Eq. 4) with off-the-shelf SDE initial-value problem (IVP) numerical solvers [32]. To obtain the numerical solution of a given SDE, we adopt the Predictor-Corrector (PC) method [23], a flexible and powerful framework for solving SDEs inherent in SGMs. The iterative steps of the Predictor-Corrector solver first use the prediction of numerical SDE solvers like the Euler–Maruyama method, and then correct the predicted trajectory with score-based MCMC methods, such as Langevin MCMC, achieving a balance of numerical accuracy and efficiency. The Algorithm  1 outlines the inference process of the denoising model, incorporating annealed Langevin dynamics as a corrector ($\epsilon_{i}$ represents the step size for Langevin Dynamics).